A rectangle is drawn so the width is 5 inches longer than the height. If the rectangle's diagonal measurement is 28 inches, find the height.
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52777)
Answer by MathTherapy(10551)

A rectangle is drawn so the width is 5 inches longer than the height. If the rectangle's diagonal measurement is 28 inches, find the height.
Let height be H
Then width = H + 5
Using the Pythagorean formula, we get: H2 + (H + 5)2 = 282
H2 + H2 + 10H + 25 = 784
2H2 + 10H + 25 - 784 = 0
2H2 + 10H - 759 = 0
To get the height, solve the above equation for H, using the Quadratic Equation formula, or by Completing the Square.
Then, DISCARD the negative value for H since a measurement CANNOT be < 0.